2. Consider the following retroreflector (a device that reflects light back to its source), which consists of two perpendicular mirrors. Prove that any ray incident on this retroreflector is reflected backward regardless of its incident angle. 3. The numerical aperture (NA) of an optical fiber is defined as NA-sin Cnax), where

Consider the following retroreflector (a device that reflects light back to its source), which consists of two perpendicular mirrors. Prove that any ray incident on this retroreflector is reflected backward regardless of its incident angle
Answer #1

92Mirror AO is perpendicular to mirror OD

i.e angle AOD= 90°

To prove any incident rays get reflected backward regardless of its angle of incidence , it is enough to prove incident rays and reflected rays become parallel to each other after reflection from retroreflector.

Proof:

Let PB is incident ray that falls at B making an angle Theta 1 with the normal BM

BC is the reflected ray that makes angle Theta 2 with normal BM

Now the reflected ray will act as incident ray for the mirror OD making an angle Theta 3 with normal CM

CQ is the reflected ray making angle Theta 4 with normal

Now,

BM is normal to mirror OA and CM is normal to OD

angle BOC = 90°

Hence closed figure BOCM is a rectangle

and angle BMC= 90°

Now coming to traingle BCM

Sum of angle of traingle is 180°

Theta2 +Theta3+ angle BMC = 180

Theta2+ Theta 3 = 180° - 90° = 90°

....... eqn1

Now multiply eqn 1 by 2

2*(Theta2 +Theta3) = 90* 2

(2 * Theta 2 + 2*Theta3 )= 180°

(Theta1+Theta2 )+ (Theta3+Theta4)= 180°

anglePMC + angle BCQ = 180°

{ Note: angle of incidence and angle of reflection are equal

hence ,angle PBM = angle CBM

and angle BCQ =angle QCM ( see diagram)}

Hence sum of interior angle on same side of transversal. ( supplementary angles )= 180°

This proves that Ray PB and CQ are parallel to each other

Conclusion :

As rays of incident and final reflected rays become parallel to each other in retroreflector, this implies whatever may be the angle of incidence is! final reflected rays will always be parallel to incidence rays regardless of its angle of incidence